Mordica Company’s standard labor cost per unit of output is $19.00 (1.90 hours x $10.00 per hour). During August, the company incurs 2,090 hours of direct labor at an hourly cost of $11.00 per hour in making 1,000 units of finished product.

Compute the total, price, and quantity labor variances. (Round answers to 2 decimal places, e.g. 52.75.)

Total labor variance

$enter a dollar amount rounded to 2 decimal places

select an option                                                          UnfavorableFavorableNeither favorable nor unfavorable
Labor price variance

$enter a dollar amount rounded to 2 decimal places

select an option                                                          Neither favorable nor unfavorableUnfavorableFavorable
Labor quantity variance

$enter a dollar amount rounded to 2 decimal places

select an option                                                          UnfavorableNeither favorable nor unfavorableFavorable
Answer #1

Total labor variance = (19*1,000) - (2090*11)

= 3990 Unfavorable

Labor price Variance = (SR-AR) *AH

= (10 - 11)*2090

= 2090 Unfavorable

Labor quantity Variance = (SH-AH) *SR

= (1.90*1,000 - 2090)*10

= 1900 Unfavorable